OpenStudy (anonymous):

HELP! identify the error made in simplifying this expression. 6p(-2p-5p)=-p

OpenStudy (anonymous):

a + sign is missing

OpenStudy (anonymous):

6p+(-2p-5p)=-p

OpenStudy (anonymous):

6p(-2p-5p)= 6p(-7p) this is where the error occurred. they added instead of multiplying

OpenStudy (anonymous):

6p+(-7p)=-p but 6p(-7p)=-42p^2

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

or as mik_wind said they missed the '+' sign both works